# Convergence to equilibrium for a degenerate three species reaction-diffusion system

### v1

@@ Abstract.
In this work, we study a reaction-diffusion system for three species. The nonlinearity in the system comes from an underlying chemical reaction. Our main objective is to understand the long time behaviour of solutions to this reaction-diffusion system when there are degeneracies. More precisely, we treat cases when one of the diffusion coefficients vanishes while the other two diffusion coefficients stay positive. We prove convergence to equilibrium type results. In all our results, the constants appearing in the decay estimates are explicit.
@@ 1. Introduction
Reaction-diffusion equations are among the most widely used differential equations in applications. These equations govern the evolution (in time) of species concentrations at various spatial locations that are simultaneously diffusing and undergoing chemical reactions. We consider a reaction-diffusion model that concerns the three species $X_{1},X_{2}$ and $X_{3}$ undergoing the following reversible reaction:
$X_{1}+X_{2}\rightleftharpoons X_{3}$
@@
The spatial domain is taken to be a bounded domain $\Omega\subset\mathbb{R}^{N}$ with $\mathrm{C}^{2+\nu}$ boundary with $\nu>0$. For the unknown $a,b,c:[0,T)\times\Omega\to\mathbb{R}$, representing the concentrations of 3 species $X_{1},X_{2},X_{3}$ respectively, the reaction-diffusion model is given by the following system of partial differential equations:
@@
(1) $\partial_{t}a-d_{a}\Delta a&=c-ab\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ \partial_{t}b-d_{b}\Delta b&=c-ab\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ \partial_{t}c-d_{c}\Delta c&=ab-c\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ n\cdot\nabla_{x}a=n\cdot\nabla_{x}b&=n\cdot\nabla_{x}c=0\quad\mbox{ on }(0,T)\times\partial\Omega,\\ a(0,\cdot)=a_{0};\,b(0,\cdot)=b_{0};\,&c(0,\cdot)=c_{0}\qquad\qquad\mbox{ in }\Omega.\end{aligned}\right.$
@@
Here $n(x)$ denotes the outward unit normal to $\Omega$ at the point $x\in\partial\Omega$. The initial data $a_{0},b_{0},c_{0}$ are taken to be smooth and the diffusion coefficients $d_{a},d_{b},d_{c}$ are taken to be nonnegative. When the diffusion coefficients are strictly positive (referred to, now onwards, as the non-degenerate setting), it is well-known that a unique strictly positive global-in-time $\mathrm{C}^{\infty}$ solution exists for the above model [Pie04, Pie10], where the positiveness comes from the quasi-positive nature of the rate function [Ama85, QS19, Rot06]. The large time behaviour of solutions to (1) in the non-degenerate setting was addressed in [DF06] using the method of entropy. In this method, a Lyapunov functional (termed entropy) is found for the evolution equation. The negative of the time derivative of this entropy functional is referred to as the entropy dissipation functional. The entropy dissipation functional is then related back to the relative entropy via a functional inequality. This will then be followed by a Grönwall type argument to deduce convergence of relative entropy to zero. A Cziszár-Kullback-Pinsker type inequality that relates relative entropy and the $\mathrm{L}^{1}$-norm helps the authors in [DF06] to prove the convergence to equilibrium in the $\mathrm{L}^{1}(\Omega)$-norm and that the decay is exponentially fast in time. In the proof of [DF06], uniform boundedness of the solution $(a,b,c)$ to (1) plays an important role. This uniform bound is available in the non-degenerate case.
@@
In this paper, we discuss the large time behaviour of solutions to couple of degenerate cases (i.e when one of the species stops diffusing) of the three species model (1). Our work is heavily inspired by [EMT20] which dealt with a four species degenerate model where one of the species stops diffusing. The authors in [EMT20] demonstrated a so-called indirect diffusion effect wherein an effective diffusion is felt by the non-diffusive species, thanks to interplay between the diffusion from diffusive species and the underlying reversible reaction. Our work demonstrates that a similar indirect diffusion effect is present in the above three species model in the presence of degeneracies. The two degenerate reaction-diffusion systems that we study in this article are:
@@
(2) $\partial_{t}a-d_{a}\Delta a&=c-ab\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ \partial_{t}b&=c-ab\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ \partial_{t}c-d_{c}\Delta c&=ab-c\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ n\cdot\nabla_{x}a=n\cdot\nabla_{x}b&=n\cdot\nabla_{x}c=0\quad\mbox{ on }(0,T)\times\partial\Omega\\ a(0,\cdot)=a_{0};\,b(0,\cdot)=b_{0};\,&c(0,\cdot)=c_{0}\qquad\qquad\mbox{ in }\Omega.\end{aligned}\right.$
@@
(3) $\partial_{t}a-d_{a}\Delta a&=c-ab\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ \partial_{t}b-d_{b}\Delta b&=c-ab\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ \partial_{t}c&=ab-c\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ n\cdot\nabla_{x}a=n\cdot\nabla_{x}b&=n\cdot\nabla_{x}c=0\quad\mbox{ on }(0,T)\times\partial\Omega\\ a(0,\cdot)=a_{0};\,b(0,\cdot)=b_{0};\,&c(0,\cdot)=c_{0}\qquad\qquad\mbox{ in }\Omega.\end{aligned}\right.$
Here, the initial data are assumed to be smooth and strictly positive. The diffusion coefficients $d_{a},d_{c}$ in (2) and the diffusion coefficients $d_{a},d_{b}$ in (3) are assumed to be strictly positive. The existence of a smooth positive solution to (2) was proved in [DF15, Theorem 3.1] for all smooth initial non-negative data. The existence of a smooth positive solution to (3) was proved in [DF15, Theorem 3.2] for all smooth initial non-negativce data if the dimension $N\leq 3$. The authors in [DF15], however, prove the existence of a weak global-in-time positive solution to (3) in any arbitrary dimension. The process of establishing various estimates in the article [EMT20] has greatly motivated us to establish similar kind of estimates in dimension larger than three for the degenerate case (2) corresponding to $d_{b}=0$. We further use various Neumann Green’s function results from [FMT19, Rot06, Mor83, CK15], to get various estimates specifically for dimensions less than four.
@@
(4) $0<M_{1}|\Omega|=:\int_{\Omega}a(t,x)+c(t,x)=\int_{\Omega}a(0,x)+c(0,x)$
(5) $0<M_{2}|\Omega|=:\int_{\Omega}b(t,x)+c(t,x)=\int_{\Omega}b(0,x)+c(0,x)$
An equilibrium state $(a_{\infty},b_{\infty},c_{\infty})$ associated with these models should also satisfy these conservation properties. Hence we should have
(6) $a_{\infty}+c_{\infty}=M_{1},\qquad b_{\infty}+c_{\infty}=M_{2}$
Moreover, at equilibrium the rate function should vanish, i.e.
(7) $c_{\infty}=a_{\infty}b_{\infty}$
@@
From now onwards, we will be considering this unique homogeneous equilibrium state $(a_{\infty},b_{\infty},c_{\infty})$.
Some Notation:
  • For any function $f:\Omega\to\mathbb{R}$ we denote the average of $f$ as
    $\overline{f}=\frac{1}{|\Omega|}\int_{\Omega}f(x)\,{\rm d}x.$
  • $\Omega_{\tau,T}=(\tau,T)\times\Omega$ and $\Omega_{t}=(0,t)\times\Omega$
  • For $a,b,c$ non-negative, we define $A:=\sqrt{a};\ B:=\sqrt{b};\ C:=\sqrt{c}$. The deviations from their averages are denoted by
    $\delta_{A}:=A-\overline{A};\quad\delta_{B}:=B-\overline{B};\quad\delta_{C}:=C-\overline{C}.$
@@
(10) $|{d_{a}-d_{c}}|<\frac{2}{C^{PRC}_{{\frac{d_{a}+d_{c}}{2}},p^{{}^{\prime}}}}\qquad\text{ and }\qquad\frac{|d_{a}-d_{c}|}{d_{a}+d_{c}}<\frac{1}{C_{SOR}(\Omega,N,p^{{}^{\prime}})},$
where the constants $C^{PRC}_{{\frac{d_{a}+d_{c}}{2}},p^{{}^{\prime}}}$ and $C_{SOR}(\Omega,N,p^{{}^{\prime}})$, are the parabolic regularity constant (see Theorem 12 in the Appendix) and the second order regularity constant (see Theorem 11 in the Appendix), respectively.
@@
Now based on solutions of two degenerate systems we define two different entropy dissipation functional which will be related to the entropy functional by a sign of time derivative.
We define the following positive entropy dissipation functionals for $d_{b}=0$ and $d_{c}=0$:
@@
(11) $\text{For}\ d_{b}&=0\\ D(a,b,c)=&d_{a}\int_{\Omega}\frac{|\nabla a|^{2}}{a}\ +d_{c}\int_{\Omega}\frac{|\nabla c|^{2}}{c}\ +\int_{\Omega}(ab-c)\ln\Big{(}\frac{ab}{c}\Big{)}.\\ \text{For}\ d_{c}&=0\\ D(a,b,c)=&d_{a}\int_{\Omega}\frac{|\nabla a|^{2}}{a}\ +d_{b}\int_{\Omega}\frac{|\nabla b|^{2}}{b}\ +\int_{\Omega}(ab-c)\ln\Big{(}\frac{ab}{c}\Big{)}.\end{aligned}\right.$
Using classical Poincaré inequality, taking $C(\Omega)$ as the Poincaré constant of the domain, and on the last term application of the following algebraic inequality
$(x-y)(\ln{x}-\ln{y})\geq 4(\sqrt{x}-\sqrt{y})^{2};\ \ \forall x,y\geq 0,\ \text{yields:}$:

### v2

@@ Abstract.
In this work, we study a $3\times 3$ triangular reaction-diffusion system. Our main objective is to understand the long time behaviour of solutions to this reaction-diffusion system when there are degeneracies. More precisely, we treat cases when one of the diffusion coefficients vanishes while the other two diffusion coefficients stay positive. We prove convergence to equilibrium type results. In all our results, the constants appearing in the decay estimates are explicit.
Key words. Reaction-Diffusion, Entropy method, Degenerate diffusion, Large time behaviour
Mathematics Subject Classification. 35K57, 35K65, 35Q92, 92E20, 39B62
@@ 1. Introduction
Reaction-diffusion equations are among the most widely used differential equations in applications. These equations govern the evolution (in time) of species concentrations at various spatial locations that are simultaneously diffusing and undergoing chemical reactions. We consider a reaction-diffusion system that concerns the three species $\mathcal{X}_{1},\mathcal{X}_{2}$ and $\mathcal{X}_{3}$ involved in the following reversible reaction:
[…]
$\mathcal{X}_{1}$ + $\mathcal{X}_{2}$ <>[ ][ ] $\mathcal{X}_{3}$
@@
The spatial domain is taken to be a bounded domain $\Omega\subset\mathbb{R}^{N}$ with $\mathrm{C}^{2+\nu}$ boundary with $\nu>0$. For the unknowns $a,b,c:[0,T)\times\Omega\to\mathbb{R}$, representing the concentrations of the three species $\mathcal{X}_{1},\mathcal{X}_{2},\mathcal{X}_{3}$ respectively, we consider the following system of differential equations:
@@
(1) $\partial_{t}a-d_{a}\Delta a&=c-ab\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ \partial_{t}b-d_{b}\Delta b&=c-ab\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ \partial_{t}c-d_{c}\Delta c&=ab-c\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ \nabla_{x}a\cdot n=\nabla_{x}b\cdot n&=\nabla_{x}c\cdot n=0\quad\mbox{ on }(0,T)\times\partial\Omega,\\ a(0,\cdot)=a_{0};\,b(0,\cdot)=b_{0};\,&c(0,\cdot)=c_{0}\qquad\qquad\mbox{ in }\Omega.\end{aligned}\right.$
@@
Here $n(x)$ denotes the outward unit normal to $\Omega$ at the point $x\in\partial\Omega$. The initial data $a_{0},b_{0},c_{0}$ are taken to be smooth up to the closure of the domain and strictly positive. The diffusion coefficients $d_{a},d_{b},d_{c}$ are taken to be nonnegative. When all the diffusion coefficients are strictly positive (referred to, now onwards, as the non-degenerate setting), it is well-known that a unique strictly positive global-in-time $\mathrm{C}^{\infty}$ solution exists for the above model (see [11, 12]), where the positiveness comes from the quasi-positive nature of the rate function [14, 1, 13]. The long time behaviour of solutions to (1) in the non-degenerate setting was addressed in [4] using the method of entropy. In this method, a Lyapunov functional (termed entropy) is found for the evolution equation. The negative of the time derivative of this entropy functional is referred to as the entropy dissipation functional. The entropy dissipation functional is then related back to the relative entropy via a functional inequality. This will then be followed by a Grönwall type argument to deduce convergence of relative entropy to zero. A Cziszár-Kullback-Pinsker type inequality that relates relative entropy and the $\mathrm{L}^{1}$-norm helps the authors in [4] to prove the convergence to equilibrium in the $\mathrm{L}^{1}(\Omega)$-norm and to deduce that the decay is exponentially fast in time. In the proof of [4], uniform boundedness of the solution $(a,b,c)$ to (1) plays an important role. This uniform bound is available in the non-degenerate case.
@@
In this paper, we discuss the long time behaviour of solutions to a couple of degenerate cases of the three species model (1). Our work is heavily inspired by [7] which dealt with a four species degenerate model where one of the species stops diffusing. The authors in [7] demonstrated a so-called indirect diffusion effect wherein an effective diffusion is felt by the non-diffusive species, thanks to the interplay between the diffusion from diffusive species and the underlying reversible reaction. Our work demonstrates that a similar indirect diffusion effect is present in the above three species model in the presence of degeneracies. The two degenerate reaction-diffusion systems that we study in this article are:
@@
(2) $\partial_{t}a-d_{a}\Delta a&=c-ab\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ \partial_{t}b&=c-ab\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ \partial_{t}c-d_{c}\Delta c&=ab-c\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ \nabla_{x}a\cdot n=\nabla_{x}b\cdot n&=\nabla_{x}c\cdot n=0\quad\mbox{ on }(0,T)\times\partial\Omega,\\ a(0,\cdot)=a_{0};\,b(0,\cdot)=b_{0};\,&c(0,\cdot)=c_{0}\qquad\qquad\mbox{ in }\Omega,\end{aligned}\right.$
@@
(3) $\partial_{t}a-d_{a}\Delta a&=c-ab\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ \partial_{t}b-d_{b}\Delta b&=c-ab\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ \partial_{t}c&=ab-c\qquad\qquad\mbox{ in }(0,T)\times\Omega,\\ \nabla_{x}a\cdot n=\nabla_{x}b\cdot n&=\nabla_{x}c\cdot n=0\quad\mbox{ on }(0,T)\times\partial\Omega,\\ a(0,\cdot)=a_{0};\,b(0,\cdot)=b_{0};\,&c(0,\cdot)=c_{0}\qquad\qquad\mbox{ in }\Omega.\end{aligned}\right.$
Here, the initial data are assumed to be smooth and strictly positive. The diffusion coefficients $d_{a},d_{c}$ in (2) and the diffusion coefficients $d_{a},d_{b}$ in (3) are assumed to be strictly positive. The existence of a smooth global-in-time positive solution to (2) was proved in [5, Theorem 3.1]. The existence of a smooth global-in-time positive solution to (3) was proved in [5, Theorem 3.2] if the dimension $N\leq 3$. The authors in [5], however, prove the existence of a weak global-in-time positive solution to (3) in any arbitrary dimension.
The process of establishing various estimates in [7] has greatly motivated us to establish similar kind of estimates in dimension larger than three for the degenerate case (2) corresponding to the vanishing of the diffusion coefficient $d_{b}$. We further use various Neumann Green’s function results from [14, 10, 3, 8] to get various estimates specifically for dimensions less than four.
@@
(4) $0<M_{1}|\Omega|=:\int_{\Omega}\big{(}a(t,x)+c(t,x)\big{)}\,{\rm d}x=\int_{\Omega}\big{(}a(0,x)+c(0,x)\big{)}\,{\rm d}x,$
(5) $0<M_{2}|\Omega|=:\int_{\Omega}\big{(}b(t,x)+c(t,x)\big{)}\,{\rm d}x=\int_{\Omega}\big{(}b(0,x)+c(0,x)\big{)}\,{\rm d}x.$
An homogenous (constant) equilibrium state $(a_{\infty},b_{\infty},c_{\infty})$ associated with these models should also satisfy the above conservation properties. Hence we should have
(6) $a_{\infty}+c_{\infty}=M_{1},\qquad b_{\infty}+c_{\infty}=M_{2}.$
Moreover, at equilibrium, the rate function should vanish, i.e.
(7) $c_{\infty}=a_{\infty}b_{\infty}.$
@@
The corresponding $a_{\infty}$ and $b_{\infty}$ can be computed using (6). Now onwards, we will be considering this unique homogeneous equilibrium state $(a_{\infty},b_{\infty},c_{\infty})$. Next, we list a bunch of notations which will be used throughout this manuscript. These notations are inspired by those used in [4].
  • The square roots of the species concentrations are denoted as
    $A:=\sqrt{a},\quad B:=\sqrt{b}=B,\quad C:=\sqrt{c}.$
  • The square roots of the homogeneous equilibrium states are denoted as
    $A_{\infty}:=\sqrt{a_{\infty}},\quad B_{\infty}:=\sqrt{b_{\infty}},\quad C_{\infty}:=\sqrt{c_{\infty}}.$
  • The average of a function $f:\Omega\to\mathbb{R}$ is denoted as
    $\overline{f}:=\frac{1}{|\Omega|}\int_{\Omega}f(x)\,{\rm d}x.$
  • The deviations of the square roots of species concentrations from their averages are denoted as
    $\delta_{A}:=A-\overline{A},\qquad\delta_{B}:=B-\overline{B},\qquad\delta_{C}:=C-\overline{C}.$
  • The parabolic cylinders are denotes as
    $\Omega_{\tau,T}:=(\tau,T)\times\Omega\qquad\mbox{ for }\,0\leq\tau<T.$
  • The lateral boundary of the parabolic cylinders are denoted as
    $\partial\Omega_{\tau,T}:=(\tau,T)\times\partial\Omega\qquad\mbox{ for }\,0\leq\tau<T.$
@@
(10) $|{d_{a}-d_{c}}|<\frac{2}{C^{PRC}_{{\frac{d_{a}+d_{c}}{2}},p^{\prime}}}\qquad\text{ and }\qquad\frac{|d_{a}-d_{c}|}{d_{a}+d_{c}}<\frac{1}{C_{SOR}(\Omega,N,p^{\prime})},$
where the constants $C^{PRC}_{{\frac{d_{a}+d_{c}}{2}},p^{\prime}}$ and $C_{SOR}(\Omega,N,p^{\prime})$, are the parabolic regularity constant (see Theorem 12 in the Appendix) and the second order regularity constant (see Theorem 11 in the Appendix), respectively.
@@
Observe that our above result is unconditional when the dimension $N\leq 3$, i.e. there is no closeness assumption on the non-zero diffusion coefficients $d_{a}$ and $d_{c}$. Furthermore, the constants $M_{1}$ and $M_{2}$ are determined by the initial data (see (4) and (5)).
@@
In the reminder of this introduction, we briefly describe our strategy to arrive at the aforementioned large time behaviour via the entropy method. We start by defining the entropy dissipation functionals associated with the degenerate models (2) and (3):
@@
(11) $\text{For}\ d_{b}&=0,\\ D(a,b,c):=&\,4\,d_{a}\int_{\Omega}\left|\nabla\sqrt{a}\right|^{2}\,{\rm d}x+\,4\,d_{c}\int_{\Omega}\left|\nabla\sqrt{c}\right|^{2}\,{\rm d}x+\int_{\Omega}(ab-c)\ln\left(\frac{ab}{c}\right)\,{\rm d}x,\\ \text{For}\ d_{c}&=0,\\ D(a,b,c):=&\,4\,d_{a}\int_{\Omega}\left|\nabla\sqrt{a}\right|^{2}\,{\rm d}x+\,4\,d_{b}\int_{\Omega}\left|\nabla\sqrt{b}\right|^{2}\,{\rm d}x+\int_{\Omega}(ab-c)\ln\left(\frac{ab}{c}\right)\,{\rm d}x.\end{aligned}\right.$
Note that both the dissipation functionals are positive. We recall here an algebraic inequality:
$(x-y)(\ln{x}-\ln{y})\geq 4(\sqrt{x}-\sqrt{y})^{2},\ \ \forall x,y\geq 0.$
This algebraic inequality gives a lower bound on the last term of the above entropy dissipation functionals whereas the classical Poincaré inequality gives a lower bound for the gradient terms. More precisely, we have